OpenStudy (anonymous):

Please help! Find the positive and negative angles of smallest magnitude that are coterminal with pi/5. The smallest positive angle is...?

OpenStudy (anonymous):

pi/5 + 2pi, or \[\frac{5\pi}{2}\]

OpenStudy (anonymous):

adding 2 pi means you go completely around the circle, back to where you were originally

OpenStudy (anonymous):

so the are COterminal

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

my brain failed for a moment

OpenStudy (anonymous):

1/5 + 2 isn't 2.5, it is 2.2

OpenStudy (anonymous):

so actually \[\frac{11\pi}{5}\]

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

Oh, it's fine! Thanks for the help!

OpenStudy (anonymous):

now for the negative angle, you would just subtract 2pi, to go back all the way around the circle again, and get \[\frac{-9\pi}{5}\]

OpenStudy (anonymous):

Awesome, thanks!